Compute the interatomic distance between two atoms inside a diatomic molecule given spectroscopic data (the rotational temperature) LaTeX: \Theta_{rot} Θ r o t

This problem has two questions. The first is for a real molecule . The second is a hypothetical molecule (made up atomic masses and LaTeX: \Theta_{rot} Θ r o t )

1) compute the internuclear separation for 35Cl2 in picometers (pm). For consistent mass data use values from:

2) Consider a hypothetical diatomic molecule where the mass of atom 1 = 45.861 Daltons and the mass of atom 2 = 21.103 Daltons. If the rotational temperature LaTeX: \Theta_{rot}= Θ r o t = 3 K, what is the internuclear distance in picometers.
Answer #1

DATE: PAGE NO 20 OR 2 Ke 2 h han h. Plancks constant Rotational Temporalu KB= Boltzman constant I Moment of Inertia (ur?) uothelial dalam molecule 0 IE (6-626X10- Ux2 Ka I 4X3-14X3-14 Y 38 X 10 - JRX3 I - 0.1344 x 10-45 kg ma u= mim2 m +m = 95-86)